# Neuromorphic Realization of Best Response in Finite-Action Games

### 03 Apr 2026 v1

@@ II-A Problem Setup
We equip $\mathcal{A}$ with a symmetric action-coupling operator $A\in\mathbb{R}^{K\times K}$, where $A_{kj}$ encodes geometry or relational affinity between actions $k$ and $j$ (e.g., spatial proximity, functional substitutability, shared resources). We define utility of agent $i$ as the projection of evidence onto the dominant eigenspace of $A$: $U_{i}(a)=\Psi_{A}\!\left(\boldsymbol{b}_{i}(a)\right)$, where $\Psi_{A}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ aggregates evidence across actions weighted by coupling strength. $U_{i}\in\mathbb{R}^{K}$ is the effective utility in the game-theoretic sense [12]: it is the quantity whose maximization defines rational action selection when the action space carries relational structure. The geometry-aware best response for agent $i$ is then
@@
where $a_{-i}\in\mathcal{A}^{N-1}$ is the action profile of all agents except $i$. A profile $a^{\star}$ is a projected Nash equilibrium if $a^{\star}_{i}\in\mathrm{BR}^{A}_{i}(a^{\star}_{-i})$ for all $i\in\mathcal{N}$. In the following, we propose neuromorphic dynamics that realize this best response.
@@ II-B Neuromorphic Decision Dynamics
We propose neuromorphic decision dynamics (NDD) for a cyclic action set $\mathcal{A}$ equipped with a circulant, symmetric action-coupling operator $A$. The action set is viewed as a discrete ring, and $A$ defines a shift-invariant interaction pattern analogous to ring-attractor networks [1] whose spectral structure we exploit in Section III. Each agent $i$ maintains an internal decision $\boldsymbol{z}_{i}(t)\in\mathbb{R}^{K}$, where each component $z_{i}^{k}$ represents preference for action $k\in\mathcal{A}$ evolving according to
@@ III-A Low-dimensional reduction
Expand $F$ using $S(u)=u+\tfrac{s_{3}}{6}u^{\circ 3}+\mathcal{O}(\|u\|^{5})$ and project onto $E_{c}$: $\mathcal{G}=\Psi_{A}L(\alpha)(\boldsymbol{v}+\boldsymbol{w})+\Psi_{A}(\boldsymbol{b})+\tfrac{s_{3}}{6}\Psi_{A}\!\big(\alpha A(\boldsymbol{v}+\boldsymbol{w})+\boldsymbol{b}\big)^{\circ 3}+\mathcal{O}_{\geq 5}$ For the linear terms: $\boldsymbol{v}\in E_{c}$ is an eigenvector of $A$ with eigenvalue $\lambda_{k_{\star}}$, so $\Psi_{A}L(\alpha)\boldsymbol{v}=\mu\boldsymbol{v}$. Since $E_{s}$ is invariant under $A$ for any symmetric circulant matrix [6] (each real Fourier plane is an eigenspace), $A\boldsymbol{w}\in E_{s}$ and thus $\Psi_{A}L(\alpha)\boldsymbol{w}=0$. The linear contribution is exactly $\mu\boldsymbol{v}+\Psi_{A}(\boldsymbol{b})$. For the cubic term, expand at $(\alpha_{c},0)$ and retain the displayed leading orders: $\big(\alpha_{c}A(\boldsymbol{v}+\boldsymbol{w})+\boldsymbol{b}\big)^{\circ 3}=(\alpha_{c}A\boldsymbol{v})^{\circ 3}+3(\alpha_{c}A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b}$. Using $\|\boldsymbol{w}\|=\mathcal{O}(\|\boldsymbol{v}\|^{3})+\mathcal{O}(\|\boldsymbol{b}\|)$, the remainder terms satisfy $(\alpha_{c}A\boldsymbol{v})^{\circ 2}\circ A\boldsymbol{w}=\mathcal{O}(\|\boldsymbol{v}\|^{5})+\mathcal{O}(\|\boldsymbol{b}\|\|\boldsymbol{v}\|^{2})$, $(\alpha_{c}A\boldsymbol{v})\circ(\alpha_{c}A\boldsymbol{w}+\boldsymbol{b})^{\circ 2}=\mathcal{O}(\|\boldsymbol{b}\|^{2}\|\boldsymbol{v}\|)+\text{h.o.t.}$, and $(\alpha_{c}A\boldsymbol{w}+\boldsymbol{b})^{\circ 3}=\mathcal{O}(\|\boldsymbol{b}\|^{3})+\text{h.o.t.}$ The projected cubic expansion contains ths displayed terms $\Psi_{A}((A\boldsymbol{v})^{\circ 3})$ and $\Psi_{A}((A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b})$, while the remaining contributions, including those involving $\boldsymbol{w}$, are absorbed into $\mathcal{R}$. Collecting terms yields (1). ∎
@@ III-B Polar Reduced Equations
Substitute $v=r\boldsymbol{e}_{\theta}$ into the reduced equation (1) from Lemma 1. Since $\boldsymbol{e}_{\theta}$ is an eigenvector of $A$ with eigenvalue $\lambda_{k_{\star}}$, we have $A\boldsymbol{v}=\lambda_{k_{\star}}r\boldsymbol{e}_{\theta}$, and thus $(A\boldsymbol{v})^{\circ 3}=\lambda_{k_{\star}}^{3}r^{3}\,\boldsymbol{e}_{\theta}^{\circ 3}$ and $(A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b}=\lambda_{k_{\star}}^{2}r^{2}\,\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b}$. For the cubic self-interaction, the identity $\Psi_{A}(\boldsymbol{e}_{\theta}^{\circ 3})=\tfrac{3}{2\sqrt{2}}\boldsymbol{e}_{\theta}$ holds for circulant $A$ [6], yielding the coefficient $\Gamma=\frac{s_{3}}{6}\alpha_{c}^{3}\lambda_{k_{\star}}^{3}\cdot\frac{3}{2}=\frac{s_{3}}{4}\alpha_{c}^{3}\lambda_{k_{\star}}^{3}$. For the mixed cubic term, decompose $\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b}$ in the Fourier basis. Using $\boldsymbol{e}_{\theta}^{\circ 2}=\frac{1}{\sqrt{2}}(\Phi_{0}+\cos 2\theta\,\Phi_{2k_{\star}}+\sin 2\theta\,\Psi_{2k_{\star}})$, the convolution with $\boldsymbol{b}$ projects onto modes $k_{\star}$ and $3k_{\star}$ under $\Psi_{A}$: $\Psi_{A}(\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b})=\frac{1}{\sqrt{2}}\Big[A_{k_{\star}}^{b}\cos(\theta-\theta_{k_{\star}}^{b})\,\boldsymbol{e}_{\theta}+A_{3k_{\star}}^{b}\cos(3\theta-\theta_{3k_{\star}}^{b})\,\boldsymbol{e}_{\theta}+\text{angular terms}\Big]$, with coefficient $\eta=\frac{s_{3}}{4}\alpha_{c}^{2}\lambda_{k_{\star}}^{2}$. Similarly, projecting $\mathcal{G}=0$ onto $\boldsymbol{e}_{\theta}$ (radial) and $\boldsymbol{e}_{\theta+\pi/2}$ (angular) yields (5)–(6) directly, while the remaining contributions, including those induced through the reduction map $\boldsymbol{w}$ are absorbed into the stated remainder terms in (1). ∎
@@ III-C Supercritical Bifurcation and Decision Formation
The range equation is $\mathcal{H}(\boldsymbol{v},\boldsymbol{w},\alpha,\boldsymbol{b})=\Psi_{A}^{\perp}F(\boldsymbol{v}+\boldsymbol{w},\alpha,\boldsymbol{b})=0$. Since $D_{w}\mathcal{H}(0,0,\alpha_{c},0)=\Psi_{A}^{\perp}L(\alpha_{c})$ is invertible on $E_{s}$, the implicit function theorem gives $\boldsymbol{w}=-\bigl(L(\alpha_{c})\bigr)^{-1}\!\left(\Psi_{A}^{\perp}\boldsymbol{b}+\mathrm{h.o.t.}\right)$. In the Fourier basis of the circulant operator $A$, this yields $w_{m}=-b_{m}/\sigma_{m}+\mathrm{h.o.t.}$ for each $m\neq k_{\star}$. Therefore, under (7), these modes are attenuated by the stable spectrum and contribute only higher-order corrections to the reduced angular equation. ∎
@@ III-E Realization of Best Response
This perspective also suggests an inverse design problem: given a desired projection of evidence onto a low-dimensional subspace, synthesize an action-coupling operator $A$ whose dominant eigenstructure realizes that projection. With graph constraints, this becomes a structured eigenstructure assignment problem. In symmetric settings (ring), the admissible eigenvectors are fixed by symmetry, and design acts through eigenvalue placement among Fourier modes. Finally, while the evidence lies in $\mathbb{R}^{K}$, the effective feature dimension is determined by the dominant eigenspace of $A$. In this sense, the proposed dynamics combine linear function approximation with a mechanistic dynamical-systems realization of best response.
@@ IV A Coverage Game Analysis
In the proposed framework, the projected utility $U_{ik}=[\Psi_{A}(\boldsymbol{b}_{i}(t))]_{k}$ is a linear projection of the evidence field, with the projection determined by the dominant eigenspace of the action-coupling operator $A$. Thus, $A$ specifies which evidence components become decision-relevant in realizing the best response. The coverage game is a natural example because the action utility is a linear functional of the underlying evidence, but the same structure also arises in other finite-action problems with relational action structure, including correlated task allocation, discretized heading selection, and structured hypothesis selection.
We consider a discrete ring of $K$ sectors $\mathcal{A}=\{1,\dots,K\}$ with event density $V:\mathcal{A}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, $\sum_{k}V(k,t)=1$, varying on a slow time scale $\tau_{b}\gg\tau_{z}$ relative to (2). Each agent observes evidence $b_{ik}(t)=V(k,t)-\rho\sum_{j\neq i}\mathbf{1}[a_{j}(t)=k]$, $\rho>0$, balancing event density against congestion [8]. For any symmetric circulant $A$ satisfying Assumption 1, the local bifurcation analysis of Section III implies, under Assumption 2, that the dynamics (2) implicitly computes the effective utility $U_{ik}(a_{-i},t)=[\Psi_{A}(\boldsymbol{b}_{i}(t))]_{k}=\sum_{s\in\mathcal{A}}C_{ks}\,b_{is}(t)$, where $C_{ks}:=\cos\!\big(k^{\star}(\theta_{s}-\theta_{k})\big)$ are the entries of the symmetric interaction kernel $C$ induced by the dominant mode, and the locally stable equilibrium of NDD selects the corresponding projected best response near bifurcation onset. These utilities are the marginal increments of the welfare function (8), so the game is an exact potential game (Lemma 3) whose maximizers are Nash equilibria reached exponentially fast at rate $\text{min}(1,2|\mu|)$, independent of $K$ (Proposition 2). Under subcritical bifurcation, amplitude and phase hysteresis ensures robustness to perturbations below $\delta^{\star}$ (Proposition 3).
@@
Since $\tau_{b}\gg\tau_{z}$, $\boldsymbol{b}_{i}(t)$ is quasi-static and (2) tracks the moving attractor adiabatically. At each $t$, we establish convergence to a best response in amplitude and angle separately, with overall rate $\min(1,2|\mu|)$: $\alpha>\alpha_{c}$ and $\Gamma<0$ give stable branch $r^{\star}=\sqrt{-\mu/\Gamma}$ with $\partial_{r}g_{r}(r^{\star})=-2|\mu|<0$ (Section III-C), so the radial component converges exponentially at rate $2|\mu|$. $A^{b}_{k_{\star}}\neq 0$ gives a unique stable angular solution $\theta^{\star}=\theta^{b}_{k_{\star}}(t)$ for any $\theta(0)$, since $\partial_{\theta}g_{\theta}(\theta^{b}_{k_{\star}})=-1$ globally at leading order, yielding exponential convergence at rate $1$. By Section III-E and Lemma 2, the attractor $(r^{\star},\theta^{\star})$ selects $a_{i}^{\star}=\mathrm{BR}_{i}^{A}(t)$, the geometry-aware best response of agent $i$ at each $t$. The overall convergence rate is $\min(1,2|\mu|)$, determined by the slower of the two modes. ∎
@@ IV-A Numerical Experiment
We simulate a repeated coverage game with $N=10$ agents on a ring of $K=18$ sectors over $T=700$ steps. The coupling matrix $A$ is sampled from a symmetric circulant Mexican-hat kernel (Figure 2(B)), with dominant eigenmode $k_{\star}=1$ (Assumption 1). This encodes the goal of positioning each agent at the event density center of mass of its neighborhood. The event density $V(k,t)$ has three peaks of unequal amplitude concentrated near sectors $\{3,\,8\text{--}10,\,14\text{--}16\}$ (Figure 2(A)), with transient ambiguity at $t\approx 280$ and a directional shift at $t\approx 430$, on timescale $\tau_{b}\gg\tau_{z}$. The global objective is to maximize weighted coverage $W(a,t)$ (Lemma 3) whose local maximizers are under Nash equilibria. Figure 3 plots the agent trajectories (white circles), the cumulative switches, and the BR fraction (fraction of agents simultaneously playing $\mathrm{BR}^{A}_{i}(t)$ at a given step) for both NDD and logit dynamics. Subcritical NDD ($\alpha=0.96<\alpha_{c}=1$) is compared against logit dynamics which is supplied $\Psi_{A}(\boldsymbol{b}_{i})$ directly. NDD recovers $\Psi_{A}(\boldsymbol{b}_{i})$ from raw evidence $\boldsymbol{b}_{i}$ alone. Figure 3 shows that NDD outperforms logit across all four metrics. NDD agents spontaneously spread into three stable clusters around the dominant peaks of $V(\cdot,t)$, maximizing $W(a,t)$ bidirectionally, while logit agents disperse incoherently across all $18$ sectors. NDD maintains BR fraction =1.0=1.0 for all $700$ steps (Proposition 2), whereas logit fluctuates between 0 and $1$ and never stabilizes. Subcritical hysteresis (Proposition 3) renders NDD immune to the ambiguity window near $t\approx 280$ which makes the logit dynamics BR fraction collapse repeatedly. At $t\approx 430$, NDD executes a single deterministic reallocation when the directional shift exceeds $\delta^{\star}$ and immediately re-commits, accumulating $O(10)$ total switches over $700$ steps versus $O(5000)$ for logit.

### 16 Apr 2026 v2

@@ II-A Problem Setup
We equip $\mathcal{A}$ with a symmetric action-coupling operator $A\in\mathbb{R}^{K\times K}$, where $A_{kj}$ encodes geometry or relational affinity between actions $k$ and $j$ (e.g., spatial proximity, functional substitutability, shared resources). We define utility of agent $i$ as the projection of evidence onto the dominant eigenspace of AA: $U_{i}(a)=\Psi_{A}\!\left(\boldsymbol{b}_{i}(a)\right)$, where $\Psi_{A}:\mathbb{R}^{K}\to\mathbb{R}^{K}$ aggregates evidence across actions weighted by coupling strength. $U_{i}\in\mathbb{R}^{K}$ is the effective utility in the game-theoretic sense [12]: it is the quantity whose maximization defines rational action selection when the action space carries relational structure. The geometry-aware best response for agent $i$ is then
@@
where $a_{-i}\in\mathcal{A}^{N-1}$ is the action profile of all agents except ii. A profile $a^{\star}$ is a projected Nash equilibrium if $a^{\star}_{i}\in\mathrm{BR}^{A}_{i}(a^{\star}_{-i})$ for all $i\in\mathcal{N}$. In the following, we propose neuromorphic dynamics that realize this best response.
@@ II-B Neuromorphic Decision Dynamics
We propose neuromorphic decision dynamics (NDD) for a cyclic action set $\mathcal{A}$ equipped with a circulant, symmetric action-coupling operator AA. The action set is viewed as a discrete ring, and AA defines a shift-invariant interaction pattern analogous to ring-attractor networks [1] whose spectral structure we exploit in Section III. Each agent $i$ maintains an internal decision $\boldsymbol{z}_{i}(t)\in\mathbb{R}^{K}$, where each component $z_{i}^{k}$ represents preference for action $k\in\mathcal{A}$ evolving according to
@@ III-A Low-dimensional reduction
Expand $F$ using $S(u)=u+\tfrac{s_{3}}{6}u^{\circ 3}+\mathcal{O}(\|u\|^{5})$ and project onto $E_{c}$: $\mathcal{G}=\Psi_{A}L(\alpha)(\boldsymbol{v}+\boldsymbol{w})+\Psi_{A}(\boldsymbol{b})+\tfrac{s_{3}}{6}\Psi_{A}\!\big(\alpha A(\boldsymbol{v}+\boldsymbol{w})+\boldsymbol{b}\big)^{\circ 3}+\mathcal{O}_{\geq 5}$ For the linear terms: $\boldsymbol{v}\in E_{c}$ is an eigenvector of AA with eigenvalue $\lambda_{k_{\star}}$, so $\Psi_{A}L(\alpha)\boldsymbol{v}=\mu\boldsymbol{v}$. Since $E_{s}$ is invariant under AA for any symmetric circulant matrix [6] (each real Fourier plane is an eigenspace), $A\boldsymbol{w}\in E_{s}$ and thus $\Psi_{A}L(\alpha)\boldsymbol{w}=0$. The linear contribution is exactly $\mu\boldsymbol{v}+\Psi_{A}(\boldsymbol{b})$. For the cubic term, expand at $(\alpha_{c},0)$ and retain the displayed leading orders: $\big(\alpha_{c}A(\boldsymbol{v}+\boldsymbol{w})+\boldsymbol{b}\big)^{\circ 3}=(\alpha_{c}A\boldsymbol{v})^{\circ 3}+3(\alpha_{c}A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b}$. Using $\|\boldsymbol{w}\|=\mathcal{O}(\|\boldsymbol{v}\|^{3})+\mathcal{O}(\|\boldsymbol{b}\|)$, the remainder terms satisfy $(\alpha_{c}A\boldsymbol{v})^{\circ 2}\circ A\boldsymbol{w}=\mathcal{O}(\|\boldsymbol{v}\|^{5})+\mathcal{O}(\|\boldsymbol{b}\|\|\boldsymbol{v}\|^{2})$, $(\alpha_{c}A\boldsymbol{v})\circ(\alpha_{c}A\boldsymbol{w}+\boldsymbol{b})^{\circ 2}=\mathcal{O}(\|\boldsymbol{b}\|^{2}\|\boldsymbol{v}\|)+\text{h.o.t.}$, and $(\alpha_{c}A\boldsymbol{w}+\boldsymbol{b})^{\circ 3}=\mathcal{O}(\|\boldsymbol{b}\|^{3})+\text{h.o.t.}$ The projected cubic expansion contains ths displayed terms $\Psi_{A}((A\boldsymbol{v})^{\circ 3})$ and $\Psi_{A}((A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b})$, while the remaining contributions, including those involving $\boldsymbol{w}$, are absorbed into $\mathcal{R}$. Collecting terms yields (1). ∎
@@ III-B Polar Reduced Equations
Substitute $v=r\boldsymbol{e}_{\theta}$ into the reduced equation (1) from Lemma 1. Since $\boldsymbol{e}_{\theta}$ is an eigenvector of AA with eigenvalue $\lambda_{k_{\star}}$, we have $A\boldsymbol{v}=\lambda_{k_{\star}}r\boldsymbol{e}_{\theta}$, and thus $(A\boldsymbol{v})^{\circ 3}=\lambda_{k_{\star}}^{3}r^{3}\,\boldsymbol{e}_{\theta}^{\circ 3}$ and $(A\boldsymbol{v})^{\circ 2}\circ\boldsymbol{b}=\lambda_{k_{\star}}^{2}r^{2}\,\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b}$. For the cubic self-interaction, the identity $\Psi_{A}(\boldsymbol{e}_{\theta}^{\circ 3})=\tfrac{3}{2\sqrt{2}}\boldsymbol{e}_{\theta}$ holds for circulant AA [6], yielding the coefficient $\Gamma=\frac{s_{3}}{6}\alpha_{c}^{3}\lambda_{k_{\star}}^{3}\cdot\frac{3}{2}=\frac{s_{3}}{4}\alpha_{c}^{3}\lambda_{k_{\star}}^{3}$. For the mixed cubic term, decompose $\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b}$ in the Fourier basis. Using $\boldsymbol{e}_{\theta}^{\circ 2}=\frac{1}{\sqrt{2}}(\Phi_{0}+\cos 2\theta\,\Phi_{2k_{\star}}+\sin 2\theta\,\Psi_{2k_{\star}})$, the convolution with $\boldsymbol{b}$ projects onto modes $k_{\star}$ and $3k_{\star}$ under $\Psi_{A}$: $\Psi_{A}(\boldsymbol{e}_{\theta}^{\circ 2}\circ\boldsymbol{b})=\frac{1}{\sqrt{2}}\Big[A_{k_{\star}}^{b}\cos(\theta-\theta_{k_{\star}}^{b})\,\boldsymbol{e}_{\theta}+A_{3k_{\star}}^{b}\cos(3\theta-\theta_{3k_{\star}}^{b})\,\boldsymbol{e}_{\theta}+\text{angular terms}\Big]$, with coefficient $\eta=\frac{s_{3}}{4}\alpha_{c}^{2}\lambda_{k_{\star}}^{2}$. Similarly, projecting $\mathcal{G}=0$ onto $\boldsymbol{e}_{\theta}$ (radial) and $\boldsymbol{e}_{\theta+\pi/2}$ (angular) yields (5)–(6) directly, while the remaining contributions, including those induced through the reduction map $\boldsymbol{w}$ are absorbed into the stated remainder terms in (1). ∎
@@ III-C Supercritical Bifurcation and Decision Formation
The range equation is $\mathcal{H}(\boldsymbol{v},\boldsymbol{w},\alpha,\boldsymbol{b})=\Psi_{A}^{\perp}F(\boldsymbol{v}+\boldsymbol{w},\alpha,\boldsymbol{b})=0$. Since $D_{w}\mathcal{H}(0,0,\alpha_{c},0)=\Psi_{A}^{\perp}L(\alpha_{c})$ is invertible on $E_{s}$, the implicit function theorem gives $\boldsymbol{w}=-\bigl(L(\alpha_{c})\bigr)^{-1}\!\left(\Psi_{A}^{\perp}\boldsymbol{b}+\mathrm{h.o.t.}\right)$. In the Fourier basis of the circulant operator AA, this yields $w_{m}=-b_{m}/\sigma_{m}+\mathrm{h.o.t.}$ for each $m\neq k_{\star}$. Therefore, under (7), these modes are attenuated by the stable spectrum and contribute only higher-order corrections to the reduced angular equation. ∎
@@ III-E Realization of Best Response
This perspective also suggests an inverse design problem: given a desired projection of evidence onto a low-dimensional subspace, synthesize an action-coupling operator AA whose dominant eigenstructure realizes that projection. With graph constraints, this becomes a structured eigenstructure assignment problem. In symmetric settings (ring), the admissible eigenvectors are fixed by symmetry, and design acts through eigenvalue placement among Fourier modes. Finally, while the evidence lies in $\mathbb{R}^{K}$, the effective feature dimension is determined by the dominant eigenspace of AA. In this sense, the proposed dynamics combine linear function approximation with a mechanistic dynamical-systems realization of best response.
@@ IV A Coverage Game Analysis
In the proposed framework, the projected utility $U_{ik}=[\Psi_{A}(\boldsymbol{b}_{i}(t))]_{k}$ is a linear projection of the evidence field, with the projection determined by the dominant eigenspace of the action-coupling operator AA. Thus, AA specifies which evidence components become decision-relevant in realizing the best response. The coverage game is a natural example because the action utility is a linear functional of the underlying evidence, but the same structure also arises in other finite-action problems with relational action structure, including correlated task allocation, discretized heading selection, and structured hypothesis selection.
We consider a discrete ring of $K$ sectors $\mathcal{A}=\{1,\dots,K\}$ with event density $V:\mathcal{A}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$, $\sum_{k}V(k,t)=1$, varying on a slow time scale $\tau_{b}\gg\tau_{z}$ relative to (2). Each agent observes evidence $b_{ik}(t)=V(k,t)-\rho\sum_{j\neq i}\mathbf{1}[a_{j}(t)=k]$, $\rho>0$, balancing event density against congestion [8]. For any symmetric circulant AA satisfying Assumption 1, the local bifurcation analysis of Section III implies, under Assumption 2, that the dynamics (2) implicitly computes the effective utility $U_{ik}(a_{-i},t)=[\Psi_{A}(\boldsymbol{b}_{i}(t))]_{k}=\sum_{s\in\mathcal{A}}C_{ks}\,b_{is}(t)$, where $C_{ks}:=\cos\!\big(k^{\star}(\theta_{s}-\theta_{k})\big)$ are the entries of the symmetric interaction kernel $C$ induced by the dominant mode, and the locally stable equilibrium of NDD selects the corresponding projected best response near bifurcation onset. These utilities are the marginal increments of the welfare function (8), so the game is an exact potential game (Lemma 3) whose maximizers are Nash equilibria reached exponentially fast at rate $\text{min}(1,2|\mu|)$, independent of $K$ (Proposition 2). Under subcritical bifurcation, amplitude and phase hysteresis ensures robustness to perturbations below $\delta^{\star}$ (Proposition 3).
@@
Since $\tau_{b}\gg\tau_{z}$, $\boldsymbol{b}_{i}(t)$ is quasi-static and (2) tracks the moving attractor adiabatically. At each $t$, we establish convergence to a best response in amplitude and angle separately, with overall rate $\min(1,2|\mu|)$: $\alpha>\alpha_{c}$ and $\Gamma<0$ give stable branch $r^{\star}=\sqrt{-\mu/\Gamma}$ with $\partial_{r}g_{r}(r^{\star})=-2|\mu|<0$ (Section III-C), so the radial component converges exponentially at rate $2|\mu|$. $A^{b}_{k_{\star}}\neq 0$ gives a unique stable angular solution $\theta^{\star}=\theta^{b}_{k_{\star}}(t)$ for any $\theta(0)$, since $\partial_{\theta}g_{\theta}(\theta^{b}_{k_{\star}})=-1$ globally at leading order, yielding exponential convergence at rate 11. By Section III-E and Lemma 2, the attractor $(r^{\star},\theta^{\star})$ selects $a_{i}^{\star}=\mathrm{BR}_{i}^{A}(t)$, the geometry-aware best response of agent $i$ at each $t$. The overall convergence rate is $\min(1,2|\mu|)$, determined by the slower of the two modes. ∎
@@ IV-A Numerical Experiment
We simulate a repeated coverage game with $N=10$ agents on a ring of $K=18$ sectors over $T=700$ steps. The coupling matrix AA is sampled from a symmetric circulant Mexican-hat kernel (Figure 2(B)), with dominant eigenmode $k_{\star}=1$ (Assumption 1). This encodes the goal of positioning each agent at the event density center of mass of its neighborhood. The event density $V(k,t)$ has three peaks of unequal amplitude concentrated near sectors $\{3,\,8\text{--}10,\,14\text{--}16\}$ (Figure 2(A)), with transient ambiguity at $t\approx 280$ and a directional shift at $t\approx 430$, on timescale $\tau_{b}\gg\tau_{z}$. The global objective is to maximize weighted coverage $W(a,t)$ (Lemma 3) whose local maximizers are under Nash equilibria. Figure 3 plots the agent trajectories (white circles), the cumulative switches, and the BR fraction (fraction of agents simultaneously playing $\mathrm{BR}^{A}_{i}(t)$ at a given step) for both NDD and logit dynamics. Subcritical NDD ($\alpha=0.96<\alpha_{c}=1$) is compared against logit dynamics which is supplied $\Psi_{A}(\boldsymbol{b}_{i})$ directly. NDD recovers $\Psi_{A}(\boldsymbol{b}_{i})$ from raw evidence $\boldsymbol{b}_{i}$ alone. Figure 3 shows that NDD outperforms logit across all four metrics. NDD agents spontaneously spread into three stable clusters around the dominant peaks of $V(\cdot,t)$, maximizing $W(a,t)$ bidirectionally, while logit agents disperse incoherently across all 1818 sectors. NDD maintains BR fraction =1.0=1.0 for all 700700 steps (Proposition 2), whereas logit fluctuates between 0 and 11 and never stabilizes. Subcritical hysteresis (Proposition 3) renders NDD immune to the ambiguity window near $t\approx 280$ which makes the logit dynamics BR fraction collapse repeatedly. At $t\approx 430$, NDD executes a single deterministic reallocation when the directional shift exceeds $\delta^{\star}$ and immediately re-commits, accumulating $O(10)$ total switches over 700700 steps versus $O(5000)$ for logit.